Moe is three times older than Curly. Two years ago he was four times older than Curly was then. What is Curly’s age at this time?

Answer by Maths68(1474)

Let
Present age of Moe = m years old
Present age of Curly = c years old

Moe is three times older than Curly
m=3c..........(1)
Two years ago
Moe was = m-2 years old
Curly was = c-2 years old
Then
Moe was four times older than Curly was then
m-2=4(c-2)
m-2=4c-8
m=4c-8+2
m=4c-6...........(2)
Put the value of m from (1) to (2)
3c=4c-6
3c-4c=-6
-c=-6
Multiply by -1 both sides
c=6
Put the value of c in (1)
m=3c
m=3*6
m=18

Present age of Moe = m = 18 years old
Present age of Curly = c = 6 years old